Need help solving...142 is what % of 460?
Answer by stanbon(75887)

Need help solving...142 is what % of 460?
------
x/100 = 142/460
---
x = 100(71/230)
---
x = 7100/230
---
x = 30.87
So 142 is 30.87% of 460
==================
Cheers,
Stan H.